# Photoemission tomography of excitons in 2D systems: momentum-space signatures of correlated electron-hole wave functions

### v1

@@ Abstract
The momentum-space signatures of excitons can be experimentally accessed through time-resolved (pump-probe) photoelectron spectroscopy. In this work, we develop a computational framework for exciton photoemission orbital tomography (exPOT) in periodic systems, enabling the simulation and interpretation of experimental observables within many-body perturbation theory. By connecting the GWGW+Bethe-Salpeter Equation (BSE) approach to photoemission tomography, our formalism captures exciton photoemission in periodic systems, explicitly incorporating photoemission matrix element effects induced by the probe pulse. The correlated nature of electrons and holes introduces distinct consequences for excitonic photoemission, including a dependence on pump pulse polarization. Using the prototypical two-dimensional material hexagonal boron nitride, we demonstrate these effects and show how our framework extends to excitons with finite center-of-mass momentum, making it well-suited to studying momentum-dark excitons. This provides valuable insights into the microscopic nature of excitonic phenomena in quantum materials.
@@ I Introduction
Excitons—bound electron-hole pairs—play a fundamental role in determining the optical and electronic response of semiconducting and insulating (two-dimensional) materials. From a theoretical perspective, describing excitonic effects requires going beyond single-particle approximations and typically involves time-dependent density functional theory (TDDFT) with suitable long-range xc-kernels, or many-body perturbation theory (MBPT) within the $GW$ plus Bethe-Salpeter equation (BSE) framework [1, 2, 3]. Excitation energies and optical spectra computed in such a way can then be compared to absorption or luminescence spectroscopy. While theory naturally provides detailed information about the quantum mechanical nature of the exciton wave function in real and momentum space, optical experiments provide only limited microscopic insights.
@@
Here, time- and angle-resolved photoemission spectroscopy (trARPES) has emerged as a promising tool to dynamically probe excited states in energy and momentum space, thus offering enhanced insight into the microscopic nature of excitons [4, 5, 6, 7, 8, 9, 10]. It is therefore highly desirable to have a theory that can connect trARPES signature with first-principles calculations. The arguably most direct way of simulating trARPES is a numerical photodetection scheme, as implemented in real-time TDDFT [11, 12], which not only allows for pump-probe simulations in linear and nonlinear regimes but also captures final-state effects in the photoemission process [13, 14]. Unfortunately, non-equilibriumical approximation for the xc-kernel in real-time TDDFT is the adiabatic local density approximation, which fails to capture excitonic effects.
@@
This is in contrast to Green’s function methods, where the inclusion of Coulomb correlation between electron and hole, in conjunction with the screened exchange interaction, successfully describes bound electron-hole states. The non-equlibrium Greens function (NEGF) method is particularly powerful for simulating excited systems through explicit time propagation, capturing the effects of many-body interactions, such as electron-electron and electron-phonon scattering, via self-energy corrections [15]. The photoemission signal is then derived from the lesser Green’s function and the matrix element [16], with the former resulting from the non-equilibrium occupation due to the pump pulse and the latter accounting for the effects of the probe pulse. However, due to the complexity of the method, trARPES simulations within the NEGF framework have so far been restricted to model systems or Wannier Hamiltonians [17, 18], often neglecting the influence of the probe pulse [19, 20].
@@
In one-step models for photoemission, such probe pulse effects are encoded in the photoemission matrix element. Matrix element effects are not only crucial for accurately describing the photoemission process itself, but also encode rich information about quantum geometrical effects [21, 22], such as the Berry curvature [23, 24] or orbital angular momentum [25, 26]. Moreover, in organic or hybrid organic/inorganic systems, matrix element effects may easily become the dominant contribution to the observed angular dependence in the photoemission signal, forming the basis of the successful photoemission orbital tomography (POT) approach [27, 28, 29].
@@
For organic molecules, POT has recently been conceptually extended to exciton photoemission orbital tomography (exPOT) [30], which establishes a direct connection between the angular distributions observed in trARPES and the real-space and momentum-space structure of excitonic wave functions. However, this earlier work was limited to non-periodic systems, such as isolated molecules, thereby restricting its applicability to condensed-phase materials and low-dimensional crystals.
@@
In the present work, we introduce exPOT for periodic systems, an ab initio scheme to simulate trARPES from first-principles. In contrast to explicit time-propagation methods, here we work in frequency space and with the exciton wave function from GWGW+BSE calculations in the linear response regime. The photoemission signal is then expressed in terms of the Fourier-transformed single-particle Bloch functions, weighted by the BSE eigenvectors and thus reflecting the entangled many-body character of electron-hole correlation. Our numerical implementation builds upon any solution of the BSE expressed in a plane wave basis set and is thus suited to treat crystalline solids and 2D semiconductors in the repeated-slab approach. We use the plane wave approximation for the final state in the photoemission process, including effects of the light polarization. In addition, our generalized theory is also able to describe excitons with finite center-of-mass momentum, allowing for the treatment of optically dark excitons—a key feature in transition metal dichalcogenides (TMDs) and other layered materials [31, 32, 33, 8]. The exPOT approach in periodic systems thus enables access to a new class of excitonic observables in photoemission, bridging the gap between theoretical many-body wave functions and experimental momentum-resolved spectra.
@@
This work is structured as follows. First we formulate the theoretical foundations in Section II, starting from Fermi’s golden rule for the photoemission process, and including the many-body excitonic effects in terms of a Dyson orbital-like construction for Bloch functions. For energy-degenerate excitons, the resulting expression for photoemission from excitons also carries an additional dependence on the pump polarization, which is discussed thereafter. To exemplifiy the method, we present results in Section III, where we apply our method to monolayer hexagonal boron nitride (hBN), a prototypical 2D semiconductor with well-characterized excitonic properties. We further show that the resulting photoelectron angular distributions depend not only on the character of the exciton but also on the polarization of the pump pulse, and discuss the implications of our theory on the rich information contained in exciton photoemission.
@@ II.1 Derivation of exPOT for periodic systems
In this section, we derive the formalism on which we base our interpretation of photoelectron spectroscopy in terms of exciton wave functions and under the following conditions. First, we assume that the optical excitation (i.e. the pump pulse) is weak, such that we stay in the linear regime and well below the exciton Mott transition threshold, and that the pump pulse is long enough to minimize energy broadening. Second, we assume that the probe pulse interacts with the system only after the pump pulse has ended. This ensures the preparation of a coherent exciton [34, 35] and avoids energy streaking effects in the trARPES spectrum induced by the pump. Third, we assume a sufficiently long probe time window to achieve adequate energy resolution in the photoelectron spectra.
Assuming a finite band gap, we model the system’s excited state using a correlated electron-hole wavefunction, expressed in terms of single-particle Bloch states as the basis set. To distinguish between electron and hole states, we write the valence states as $\phi_{v,\boldsymbol{q}}(\boldsymbol{r})$, with $v$ denoting the band index and $\boldsymbol{q}$ the crystal momentum. Similarly, conduction states are written as $\xi_{c,\boldsymbol{q}}(\boldsymbol{r})$. With these single particle basis sets and assuming the Tamm-Dancoff approximation (TDA), we can write the $m$-th exciton wave function, $|m,\boldsymbol{Q}\rangle$, in terms of electron- and hole coordinates, $\boldsymbol{r}_{e}$ and $\boldsymbol{r}_{h}$ respectively [2]:
$\langle\boldsymbol{r}|m,\boldsymbol{Q}\rangle=$ $\Psi^{(m,\boldsymbol{Q})}(\boldsymbol{r}_{h},\boldsymbol{r}_{e})=$ (1)
$\sum_{\boldsymbol{q}}^{\mathrm{BZ}}\sum_{v}^{N_{\mathrm{occ.}}}\sum_{c}^{\infty}X_{v,c,\boldsymbol{q}}^{(m,\boldsymbol{Q})}\phi_{v,\boldsymbol{q}}^{*}(\boldsymbol{r}_{h})\xi_{c,\boldsymbol{q}+\boldsymbol{Q}}(\boldsymbol{r}_{e}).$
In this expression, $\boldsymbol{Q}$ represents the center-of-mass momentum of the exciton, and $X_{v,c,\boldsymbol{q}}^{(m,\boldsymbol{Q})}$ are the elements of the transition density matrix, which encode the exciton character and oscillator strength. These eigenstates arise from the BSE Hamiltonian or the Casida equation. The $m$-th excitation energy is further written as $\Omega^{(m,\boldsymbol{Q})}$. In the language of second quantization, such exciton states may also be written as
@@
In the spirit of ground state POT [29], and under the conditions states above, we treat the photoemission due to the probe pulse in first-order perturbation theory. Specifically, we apply Fermi’s golden rule for the transition rate from an $N$-electron initial state $|\Psi_{\mathrm{i}}^{N}\rangle$ to a final state $|\Psi_{\mathrm{f}}^{N}\rangle$:
@@
with the Dirac delta function enforcing energy conservation between the initial and final state and the energy $\omega$ carried by the photon field $\boldsymbol{A}$. Here and in the following, we use the velocity gauge in the dipole approximation and use atomic units unless denoted otherwise.
The final state $|\Psi_{\mathrm{f}}^{N}\rangle$ describes the system after the photoemission event. Here, we make use of the sudden approximation [36] and thus neglect correlation between the photoelectron described by the one-electron state $\gamma_{\boldsymbol{k}}$ and the remaining N1N-1-electron state $|\Psi_{\mathrm{f}}^{N-1}\rangle$. Therefore, $|\Psi_{\mathrm{f}}^{N}\rangle$ can be written as an anti-symmetrized product:
@@
assuming the N1N-1-electron final state to be of the form
@@
where $\hat{a}_{j,\boldsymbol{\tilde{q}}}$ annihilates an electron (creates a photo-hole) in band $j$, and with crystal momentum $\boldsymbol{\tilde{q}}$, from the $N$-electron ground state $|\Psi_{0}^{N}\rangle$.
In the following, we take the $m$-th excited state with center-of-mass momentum $\boldsymbol{Q}$ as initial state, and assume that in the final state, the emitted electron has momentum $\boldsymbol{k}$, while the photo-hole resides in $|\phi_{j,\boldsymbol{\tilde{q}}}\rangle$ (or $|\xi_{j,\boldsymbol{\tilde{q}}}\rangle$, if it were to reside in a conduction band.). Then, the matrix element $\left\langle\Psi_{\mathrm{f}}^{N}\right|\boldsymbol{A}\boldsymbol{\hat{P}}\left|\Psi_{\mathrm{i}}^{N}\right\rangle$ of Equation 3 takes the form
@@
Note that with this matrix element we can calculate the transition rate from $|\Psi_{\mathrm{m,\boldsymbol{Q}}}^{N}\rangle$ to one specific final state, $|\mathcal{A}\Psi_{j,\boldsymbol{\tilde{q}}}^{N-1}\gamma_{\boldsymbol{k}}\rangle$ where an electron has been removed from the $j$-th band (and carrying crystal momentum $\boldsymbol{\tilde{q}}$). To account for all possible photoemission final states, we will have to sum up the transition rates of all final hole quantum numbers at a later stage.
To proceed, we reduce the $N$-electron matrix element of Equation 6 to an effective one-electron matrix element by utilizing the Dyson orbital as the overlap between the $N$-electron and the N1N-1-electron system [37, 38, 39]. In contrast to existing literature for isolated systems, we here perform this step for many-body Bloch functions that carry an additional crystal momentum quantum number. To see that this is indeed possible, one can always transform Bloch functions with a finite sampling of the Brillouin-zone to a supercell. We then get for the absolute of the matrix element
@@
using the effective single-particle Bloch-Dyson state:
@@
Here the volume of integration, $V$, is the volume of the unit cell, for which we assume a sampling of the Brillouin zone in the periodic direction only, as well as sufficient vacuum in the non-periodic direction of the slab.
Next we can evaluate the formal definition of the Dyson wave function in the basis of valence- and conduction bands (see Appendix A for a derivation of this expansion):
@@
$+\sum_{c^{\prime}}\sum_{\boldsymbol{q}^{\prime}}^{\mathrm{BZ}}\langle\Psi_{j,\boldsymbol{\tilde{q}}}^{N-1}|\hat{a}_{c^{\prime},\boldsymbol{q}^{\prime}}|\Psi_{m,\boldsymbol{Q}}^{N}\rangle\xi_{c^{\prime},\boldsymbol{q}^{\prime}}(\boldsymbol{r})$ . (9)
Inserting Equation 2 and Equation 5, we get:
@@
For simplicity, we assumed the ground-state, $|\Psi_{0}^{N}\rangle$, to be a single Slater-determinant constructed from the valence states $\phi_{v,\boldsymbol{q}}$, thereby neglecting the changes in the N1N-1 electron system upon photoemission. This allows us to work with the GWGW+BSE method carried out on top of density functional theory (DFT) calculations, but we stress that our method is also suitable for more general, correlated initial states. The first term in Equation 10 containing the sum over $v^{\prime}$ then vanishes due to orthogonality of the Bloch states. By the same argument, we can set $c^{\prime}=c$, $v=j$, $\boldsymbol{q}=\boldsymbol{\tilde{q}}$ and $\boldsymbol{q}^{\prime}=\boldsymbol{q}+\boldsymbol{Q}$ in the remaining term and thus find
@@
Inserting Equation 11 into Equation 7, we get:
$\left|M^{\mathrm{(m,\boldsymbol{Q})}}_{v,\boldsymbol{q},\boldsymbol{k}}(\boldsymbol{A})\right|=\left|\int_{V}\mathrm{d}^{3}r\gamma_{\boldsymbol{k}}^{*}(\boldsymbol{r})\boldsymbol{A}\boldsymbol{\nabla}\sum_{c}X_{v,c,\boldsymbol{q}}^{(m,\boldsymbol{Q})}\xi_{c,\boldsymbol{q}+\boldsymbol{Q}}(\boldsymbol{r})\right|$ (12)
For the photoelectron wave function we follow the commonly applied strategy in POT, namely the plane wave final state (PWFS) approximation: $\gamma_{\boldsymbol{k}}(\boldsymbol{r})=e^{\mathrm{i}\boldsymbol{k}\boldsymbol{r}}/\sqrt{V}$. Integrating by parts, and dropping the surface term, then allows us to write the matrix element as a sum over Fourier-transformed Bloch functions:
@@
Before putting everything together, we have to take the energy conservation of Equation 3 into account. To this end, we identify the energy of the initial state as the sum of the ground state energy, $E_{0}^{N}$, plus the excitation energy $\Omega^{(m,\boldsymbol{Q})}$. The energy of the final state consists of the photoelectron kinetic energy, $E_{\mathrm{kin}}=\boldsymbol{k}^{2}/2$ and the energy of the N1N-1-electron state, $E_{v,\boldsymbol{q}}^{N-1}$. The latter can be expressed as the ground-state energy plus the ionization energy of the $v$-th valence band at $\boldsymbol{q}$: $\varepsilon_{v,\boldsymbol{q}}$. We then obtain
@@
Combining this expression for the energy conservation with the matrix element of Equation 13, we can now rewrite the transition rate in Equation 3. This rate gives the probability for the $m$-th exciton, with excitation energy $\Omega^{(m,\boldsymbol{Q})}$ and center-of-mass momentum $\boldsymbol{Q}$, to end up in a state with photoelectron momentum $\boldsymbol{k}$, if the photohole resides in valence state $\phi_{v,\boldsymbol{q}}$:
@@
$\cdot\delta\left(\omega+\Omega^{(m,\boldsymbol{Q})}-\varepsilon_{v,\boldsymbol{q}}-E_{\mathrm{kin}}\right).$ (15)
Summing up the transition rates for all final-state quantum-numbers $v$ and $\boldsymbol{q}$, we get an expression that is proportional to the photoelectron intensity,
@@
where $I_{m,\boldsymbol{Q}}(\boldsymbol{k})$ is the experimentally accessible number of photoelectrons of momentum $\boldsymbol{k}$, that are emitted from the $m$-th excited state carrying center-of-mass momentum $\boldsymbol{Q}$.
In conclusion, photoemission from a static exciton state is given by a sum over crystal momentum and hole state contributions, for which the photoelectron kinetic energy is given by the energy conservation between the probe-, pump-, and hole energy. At each allowed kinetic energy, the photoelectron angular distribution is given by the absolute squared of a coherent sum over electron contributions, weighted by the respective electron-hole transition coefficients that we obtain from the solutions of the BSE. In contrast to other formulations of exciton photoemission considering that only consider the spectral function, our formalism includes photoelectron matrix element effects due to the probe pulse and at the level of the PWFS approximation.

### v2

@@ Abstract
The momentum-space signatures of excitons can be experimentally accessed through time-resolved (pump-probe) photoelectron spectroscopy. In this work, we develop a computational framework for exciton photoemission orbital tomography (exPOT) in periodic systems, enabling the simulation and interpretation of experimental observables within many-body perturbation theory. By connecting the GWGW+Bethe-Salpeter equation (BSE) approach to photoemission tomography, our formalism captures exciton photoemission in periodic systems, explicitly incorporating photoemission matrix element effects induced by the light-matter interaction via the probe pulse. The correlated nature of electrons and holes introduces distinct consequences for excitonic photoemission. Using the prototypical two-dimensional material hexagonal boron nitride, we demonstrate these effects, including a dependence of the photoemission angular distribution on the pump pulse polarization. Moreover, our framework extends to excitons with finite center-of-mass momentum, making it well-suited to studying momentum-dark excitons. This provides valuable insights into the microscopic nature of excitonic phenomena in quantum materials.
@@ I Introduction
Excitons—bound electron-hole pairs—play a fundamental role in determining the optical and electronic response of semiconducting and insulating (two-dimensional) materials. From a theoretical perspective, describing excitonic effects requires going beyond single-particle approximations and typically involves time-dependent density functional theory (TDDFT) with suitable long-range xc-kernels, or many-body perturbation theory (MBPT) within the $GW$ plus Bethe-Salpeter equation (BSE) framework [1, 2, 3]. Excitation energies and optical spectra computed in such a way can then be compared to absorption or luminescence spectroscopy. While theory naturally provides detailed information about the quantum mechanical nature of the underlying exciton wave function in real and momentum space, optical experiments provide only limited microscopic insights.
@@
Here, time- and angle-resolved photoemission spectroscopy (trARPES) has emerged as a promising tool to dynamically probe excited states in energy and momentum space, thus offering enhanced insight into the microscopic nature of excitons [4, 5, 6, 7, 8, 9, 10]. It is therefore highly desirable to have a theory that can connect trARPES signatures with first-principles calculations. The arguably most direct way of simulating trARPES is a numerical photodetection scheme, as implemented in real-time TDDFT [11, 12], which not only allows for pump-probe simulations in linear and nonlinear regimes but also captures final-state effects in the photoemission process [13, 14]. Unfortunately, the only practically applicable approximation for the xc-kernel in such real-time TDDFT simulations is the adiabatic local density approximation, which fails to capture excitonic effects.
@@
This is in contrast to Green’s function methods, where the inclusion of Coulomb correlation between electron and hole, in conjunction with the screened exchange interaction, successfully describes bound electron-hole states. The non-equilibrium Greens function (NEGF) method is particularly powerful for simulating excited systems through explicit time propagation, capturing the effects of many-body interactions, such as electron-electron and electron-phonon scattering, via self-energy corrections [15]. The photoemission signal is then derived from the lesser Green’s function and the matrix element [16], with the former resulting from the non-equilibrium occupation due to the pump pulse and the latter accounting for the effects of the probe pulse. However, due to the complexity of the method, trARPES simulations within the NEGF framework have so far been restricted to model systems or Wannier-Hamiltonians [17, 18], often neglecting the transition probability to the photoelectron state induced by the probe pulse [19, 20].
@@
In one-step models for photoemission, these effects are encoded in the photoemission matrix element. Matrix element effects are not only crucial for accurately describing the photoemission process itself, but also encode rich information about quantum geometrical effects [21, 22], such as the Berry curvature [23, 24] or orbital angular momentum [25, 26]. Moreover, in organic or hybrid organic/inorganic systems, matrix element effects may easily become the dominant contribution to the observed angular dependence in the photoemission signal, forming the basis of the successful photoemission orbital tomography (POT) approach [27, 28, 29].
@@
For organic molecules, POT has recently been conceptually extended to exciton photoemission orbital tomography (exPOT) [30], which establishes a direct connection between the angular distributions observed in trARPES and the real-space and momentum-space structure of excitonic wave functions. However, this earlier work was limited to non-periodic systems, such as isolated molecules, thereby excluding its applicability to condensed-phase materials and low-dimensional crystals.
@@
In the present work, we introduce exPOT for periodic systems, an ab initio scheme to simulate trARPES from first-principles. In contrast to explicit time-propagation methods, here we work in frequency space and with the exciton wave function from GWGW+BSE calculations in the linear response regime. The photoemission signal is then expressed in terms of the Fourier-transformed single-particle Bloch functions, weighted by the BSE eigenvectors and thus reflecting the entangled many-body character of electron-hole correlation. Our numerical implementation (see Appendix B) builds upon solutions of the BSE expressed in a plane wave basis set and is thus suited to treat crystalline solids and 2D semiconductors in the repeated-slab approach. We use the plane wave approximation for the final state in the photoemission process, including effects of the light polarization. We therefore expect our approach to be applicable for probe fields in the linear response regime and for probe energies in the 15–50 eV range. Note that our generalized theory is also able to describe excitons with finite center-of-mass momentum, allowing for the treatment of optically dark excitons—a key feature in transition metal dichalcogenides and other layered materials [31, 32, 33, 8]. The exPOT approach in periodic systems thus enables access to a new class of excitonic observables in photoemission [34], bridging the gap between theoretical many-body wave functions and experimental momentum-resolved spectra.
@@
This work is structured as follows. First we formulate the theoretical foundations in Sec. II, starting from Fermi’s golden rule for the photoemission process, and including the many-body excitonic effects in terms of a Dyson orbital-like construction for Bloch functions. For energy-degenerate excitons, the resulting expression for photoemission from excitons also carries an additional dependence on the pump polarization, which is discussed thereafter. As an exemplification, we present results in Sec. III, where we apply our method to monolayer hexagonal boron nitride (hBN), a prototypical 2D semiconductor with well-characterized excitonic properties. We further show that the resulting photoelectron angular distribution maps depend not only on the character of the exciton but also on the polarization of the pump pulse, and discuss the implications of our theory on the rich information contained in exciton photoemission.
@@ II.1 Derivation of exPOT for periodic systems
In this section, we derive the formalism on which we base our interpretation of photoelectron spectroscopy in terms of exciton wave functions and under the following conditions. First, we assume that the optical excitation (i.e. the pump pulse) is weak, such that we stay in the linear regime and well below the exciton Mott transition threshold. Moreover, the pump pulse is considered to be long enough to minimize energy broadening. Second, we assume that the probe pulse interacts with the system only after the pump pulse has ended. This ensures the preparation of a relaxed exciton [35, 36] and avoids energy streaking effects in the trARPES spectrum induced by the pump field.
We consider systems with a finite band gap and model the system’s excited state using a correlated electron-hole wave function, expressed in terms of single-particle Bloch states as the basis set. To distinguish between electron and hole states, we write the valence state wave functions as $\phi_{v,\boldsymbol{q}}(\boldsymbol{r})$, with $v$ denoting the band index and $\boldsymbol{q}$ the crystal momentum. Similarly, conduction state wave functions are written as $\xi_{c,\boldsymbol{q}}(\boldsymbol{r})$. With these single particle basis sets and employing the Tamm-Dancoff approximation (TDA), we can write the $m$-th exciton wave function, $\Psi^{(m,\boldsymbol{Q})}(\boldsymbol{r}_{h},\boldsymbol{r}_{e})$, in terms of electron- and hole coordinates, $\boldsymbol{r}_{e}$ and $\boldsymbol{r}_{h}$ respectively [2]:
$\Psi^{(m,\boldsymbol{Q})}(\boldsymbol{r}_{h},\boldsymbol{r}_{e})=\sum_{\boldsymbol{q}}^{\mathrm{BZ}}\sum_{v}^{N_{\mathrm{occ.}}}\sum_{c}^{\infty}X_{v,c,\boldsymbol{q}}^{(m,\boldsymbol{Q})}\phi_{v,\boldsymbol{q}}^{*}(\boldsymbol{r}_{h})\xi_{c,\boldsymbol{q}+\boldsymbol{Q}}(\boldsymbol{r}_{e})$ (1)
In this expression, $\boldsymbol{Q}$ represents the center-of-mass momentum of the exciton, and $X_{v,c,\boldsymbol{q}}^{(m,\boldsymbol{Q})}$ encode the exciton character. They are the elements of the transition density matrix, which we obtain as eigenstates from the BSE Hamiltonian, with the corresponding excitation energies denoted as $\Omega^{(m,\boldsymbol{Q})}$. In the language of second quantization, the excitonic state may also be written as
@@
In analogy to ground state POT [29], and under the conditions stated above, we treat the photoemission due to the probe pulse in first-order perturbation theory. As outlined in Ref. [37], we apply Fermi’s golden rule for the transition rate from an $N$-electron initial state $|\Psi_{\mathrm{i}}^{N}\rangle$ to a final state $|\Psi_{\mathrm{f}}^{N}\rangle$:
@@
with the Dirac delta function enforcing energy conservation between the initial and final state and the energy $\omega$ carried by the photon field $\boldsymbol{A}$. Here and in the following, we use the velocity gauge within the dipole approximation and use atomic units unless denoted otherwise.
The final state $|\Psi_{\mathrm{f}}^{N}\rangle$ describes the system after the photoemission event. Here, we make use of the sudden approximation [38] and thus neglect correlations between the photoelectron described by the one-electron state $|\gamma_{\boldsymbol{k}}\rangle$ and the remaining (N1)(N-1)-electron state $|\Psi_{\mathrm{f}}^{N-1}\rangle$. Therefore, $|\Psi_{\mathrm{f}}^{N}\rangle$ can be written as an anti-symmetrized product:
@@
where we additionally assume the (N1)(N-1)-electron final state to be of the form
@@
with $\hat{a}_{j,\boldsymbol{\tilde{q}}}$ annihilating an electron (creating a photo-hole) in band $j$, and with crystal momentum $\boldsymbol{\tilde{q}}$, from the $N$-electron ground state $|\Psi_{0}^{N}\rangle$. Via this approximation, we neglect possible relaxations in the final (N1)(N-1)-electron system.
In the following, we take the $m$-th excited state with center-of-mass momentum $\boldsymbol{Q}$ as defined in Eq. 2 as initial state, and assume that in the final state, the emitted electron has momentum $\boldsymbol{k}$, while the photo-hole resides in $|\phi_{j,\boldsymbol{\tilde{q}}}\rangle$. Then, the matrix element $\left\langle\Psi_{\mathrm{f}}^{N}\right|\boldsymbol{A}\boldsymbol{\hat{P}}\left|\Psi_{\mathrm{i}}^{N}\right\rangle$ of Eq. 3 takes the form
@@
Note that with this matrix element, we can calculate the transition rate from $|\Psi_{\mathrm{m,\boldsymbol{Q}}}^{N}\rangle$ to one specific final state, $|\mathcal{A}\Psi_{j,\boldsymbol{\tilde{q}}}^{N-1}\gamma_{\boldsymbol{k}}\rangle$ where an electron has been removed from the $j$-th band (and carrying crystal momentum $\boldsymbol{\tilde{q}}$). To account for all possible photoemission final states, we will have to sum up the transition rates of all final hole quantum numbers at a later stage.
To proceed, we reduce the $N$-electron matrix element of Eq. 6 to an effective one-electron matrix element by introducing the Dyson orbital as the overlap between the initial $N$-electron and the final (N1)(N-1)-electron system [37, 39, 40]. In contrast to existing literature for isolated systems, we here perform this step for Bloch functions that carry an additional crystal momentum quantum number. We then get for the absolute of the matrix element
@@
where we have defined the effective single-particle Dyson wave function with Bloch momentum $\boldsymbol{\tilde{q}}$ as
@@
Here the volume of integration, $V$, is the volume of the crystal, as defined by the unit cell and the sampling of the Brillouin-zone in the periodic directions (using sufficient vacuum in the non-periodic direction of the slab).
Next we evaluate the formal definition of the Dyson wave function, starting from an alternative form, which employs the basis of valence- and conduction bands and is derived in Appendix A:
@@
$+\sum_{c^{\prime}}\sum_{\boldsymbol{q}^{\prime}}^{\mathrm{BZ}}\langle\Psi_{j,\boldsymbol{\tilde{q}}}^{N-1}|\hat{a}_{c^{\prime},\boldsymbol{q}^{\prime}}|\Psi_{m,\boldsymbol{Q}}^{N}\rangle\xi_{c^{\prime},\boldsymbol{q}^{\prime}}(\boldsymbol{r})$ (9)
Inserting Eqs. 2 and 5, we get
@@
For simplicity, we assume the ground state, $|\Psi_{0}^{N}\rangle$, to be a single Slater-determinant constructed from the Kohn-Sham valence orbitals $\phi_{v,\boldsymbol{q}}$. This allows us to work with the GWGW+BSE method carried out on top of density functional theory (DFT) calculations, but we stress that our method is also suitable for more general, correlated initial states. The first term in Eq. 10 containing the sum over $v^{\prime}$ then vanishes due to orthogonality of the Bloch states. By the same argument, we can set $c^{\prime}=c$, $v=j$, $\boldsymbol{q}=\boldsymbol{\tilde{q}}$ and $\boldsymbol{q}^{\prime}=\boldsymbol{q}+\boldsymbol{Q}$ in the second term and thus find
@@
Inserting Eq. 11 into Eq. 7, we arrive at
$\left|M^{\mathrm{(m,\boldsymbol{Q})}}_{v,\boldsymbol{q},\boldsymbol{k}}(\boldsymbol{A})\right|=\left|\int_{V}\mathrm{d}^{3}r\gamma_{\boldsymbol{k}}^{*}(\boldsymbol{r})\boldsymbol{A}\boldsymbol{\nabla}\sum_{c}X_{v,c,\boldsymbol{q}}^{(m,\boldsymbol{Q})}\xi_{c,\boldsymbol{q}+\boldsymbol{Q}}(\boldsymbol{r})\right|.$ (12)
To further evaluate the matrix element, we follow the commonly applied strategy in POT and approximate the photoelectron wave function by a plane wave final state (PWFS): $\gamma_{\boldsymbol{k}}(\boldsymbol{r})=e^{\mathrm{i}\boldsymbol{k}\boldsymbol{r}}/\sqrt{V}$. Integrating by parts, and dropping the surface term, then allows us to write the matrix element as a sum over Fourier-transformed Bloch functions:
@@
Before putting everything together, we also take the energy conservation defined in Eq. 3 into account. To this end, we identify the energy of the initial state as the sum of the ground state energy, $E_{0}^{N}$, and the excitation energy $\Omega^{(m,\boldsymbol{Q})}$. The energy of the final state consists of the photoelectron kinetic energy, $E_{\mathrm{kin}}=\boldsymbol{k}^{2}/2$ and the energy of the (N1)(N-1)-electron state, $E_{v,\boldsymbol{q}}^{N-1}$. The latter can be expressed as the ground-state energy plus the ionization energy of the $v$-th valence band at $\boldsymbol{q}$: $\varepsilon_{v,\boldsymbol{q}}$. This leads to
@@
Combining this expression for the energy conservation with the matrix element of Eq. 13, we can now rewrite the transition rate $W$ in Eq. 3. This rate gives the probability for the $m$-th exciton, with excitation energy $\Omega^{(m,\boldsymbol{Q})}$ and center-of-mass momentum $\boldsymbol{Q}$, to end up in a state with photoelectron momentum $\boldsymbol{k}$, if the photohole resides in valence orbital $\phi_{v,\boldsymbol{q}}$:
@@
$\cdot\delta\left(\omega+\Omega^{(m,\boldsymbol{Q})}-\varepsilon_{v,\boldsymbol{q}}-E_{\mathrm{kin}}\right)$ (15)
Summing up the transition rates for all final-state quantum-numbers $v$ and $\boldsymbol{q}$, we get an expression which is proportional to the overall photoelectron intensity,
@@
where $I_{m,\boldsymbol{Q}}(\boldsymbol{k})$ denotes the number of photoelectrons of momentum $\boldsymbol{k}$, that are emitted from the $m$-th excited state carrying center-of-mass momentum $\boldsymbol{Q}$.
In summary, photoemission from a static (or, in practice, long-lived, quasi-static) exciton state is given by a sum over crystal momentum and hole state contributions, for which the photoelectron kinetic energy is determined by the energy conservation between the probe-, pump-, and hole energy. At each allowed kinetic energy, the photoelectron angular distribution is given by the absolute squared of a coherent sum over electron contributions, weighted by the respective electron-hole transition coefficients, which are obtained from the solutions of the BSE. In contrast to other formulations of exciton photoemission that only consider the spectral function [19, 20], our formalism includes photoelectron matrix element effects due to the probe pulse and at the level of the PWFS approximation.